OpenStudy (anonymous):

help: This one is bugging the hell out of me :) For f(x)=2x^3+3x^2-36x (a) find the intervals on which f is increasing or decreasing (b) find the local maximum and minimum values of f (c) find the intervals of concavity and the inflection points

OpenStudy (anonymous):

\[f'(x)=6x^2+6x-36=6(x^2+x-6)=6(x+3)(x-2)\] zeros of the derivative (critical points) are -3 and 2 so it will be increasing on \[(-\infty, -3)\cup (2,\infty)\] and decreasing on \[(-3,2)\]

OpenStudy (anonymous):

-3 gives a local max, 2 gives a local min, and you and find concavity by the second derivative which is \[f''(x)=12x+6\] which is 0 at \[x=-\frac{1}{2}\]

OpenStudy (anonymous):

Thanks satellite73